OpenStudy (anonymous):

1. Select the phrase that correctly completes the sentence. The value of the dependent variable in a function is always __________ the value of the independent variable. (1 point) A.greater than B.dependent on C.equal to D.less than 2. Which set of output values correctly completes the function table? (1 point) A.8, 2, –12 B.8, 0, 12 C.8, 0, –12 D.–8, 0 12 @random231

OpenStudy (anonymous):

attachment goes with #2

random231 (random231):

#2 isC

OpenStudy (anonymous):

Can i ask a couple more?

random231 (random231):

sure

OpenStudy (anonymous):

#1 is self explanatory, dependent variable means to be dependent on the another variable which would be the independent variable

OpenStudy (anonymous):

3. Which function table is correct for the rule y = x – 18?

OpenStudy (anonymous):

4. Identify the rule for the function table. Input (x) Output (y) 24 12 12 6 –12 –6 (1 point) A.y = 12 – x B.y = x – 12 C.y = x ÷ 2 D.y = 2 ÷ x

OpenStudy (anonymous):

@Johnbc

OpenStudy (jlg030597):

How can i help you?

OpenStudy (anonymous):

When you put x = 24 into your function you get y = 12 which is half When you put x = 12 into your function you get y = 6 which is half When you put x = -12 into your function you get y = -6 which is half So what must be your function?

OpenStudy (anonymous):

For the tables just plug in the x into your y = function and see what number y would be if you plugged that number in and you will see the relationship

OpenStudy (anonymous):

wait so for #3 i think the answer is A

OpenStudy (anonymous):

For\[y=x-18\] if i plug in x to be 5 then my y would be 5-18 = -13 = y so its just a matter of trial and error

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Thx im gonna try to find #4 hold o tell me if i am right :)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

The answer for #4 is C

random231 (random231):

yup!

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Great job